An angle is 3 degrees more than twice its supplement. Find the measure of each angle.
Answer by arKed(27)

n1 + n2 = 180 eq. 1
n2 = 3 + 2(n1)
sol'n:
n1 + 3 + 2 (n1) = 180
3 n1 = 180 - 3
3 n1 / 3= 177 / 3
n1 = 59 degrees ans.
n2 = 3 + 2(n1)
n2 = 3 + 2 (59)
n2 = 121 degrees ans.